### v1

# MirrorGuard: Adaptive Defense Against Jailbreaks via Entropy-Guided Mirror Crafting

Abstract
Defending large language models (LLMs) against jailbreak attacks is crucial for ensuring their safe deployment. Existing defense strategies generally rely on predefined static criteria to differentiate between harmful and benign prompts. However, such rigid rules are incapable of accommodating the inherent complexity and dynamic nature of real jailbreak attacks. In this paper, we propose a novel concept of “mirror” to enable dynamic and adaptive defense. A mirror refers to a dynamically generated prompt that mirrors the syntactic structure of the input while ensuring semantic safety. The personalized discrepancies between the input prompts and their corresponding mirrors serve as the guiding principles for defense. A new defense paradigm, MirrorGuard, is further proposed to detect and calibrate risky inputs based on such mirrors. An entropy-based detection metric, Relative Input Uncertainty (RIU), is integrated into MirrorGuard to quantify the discrepancies between input prompts and mirrors. MirrorGuard is evaluated on several popular datasets, demonstrating state-of-the-art defense performance while maintaining general effectiveness.
MirrorGuard: Adaptive Defense Against Jailbreaks via Entropy-Guided Mirror Crafting
Rui Pu1, Chaozhuo Li1, Rui Ha1, Litian Zhang2, Lirong Qiu1, Xi Zhang1 1Beijing University of Posts and Telecommunications, 2Beihang University, {puruirui, lichaozhuo, harry, qiulirong, zhangx}@bupt.edu.cn, litianzhang@buaa.edu.cn
1 Introduction
Securing LLMs has emerged as a critical challenge due to their widespread deployment in various essential domains (OpenAI, 2023; Zhu et al., 2024). Operating within open scenarios, LLMs are inherently susceptible to jailbreak attacks, which exploit vulnerabilities to circumvent LLM safety protocols, thereby compelling models to generate harmful outputs (Liu et al., 2024; Chen et al., 2024).
Defense strategies against jailbreak attacks can be broadly categorized into prefill-level (Xie et al., 2023; Phute et al., 2023) and response-level approaches (Phute et al., 2023). Prefill-level defenses enhance the robustness of initial prompts or input structures, using methods like protective prefixes and perplexity-based classification to filter malicious inputs (Zheng et al., 2024; Alon and Kamfonas, 2023). In contrast, response-level strategies focus on evaluating model-generated outputs, applying safety filters to mitigate potential harm (Phute et al., 2023; Xu et al., 2024).
Figure 1: The differences between static discrimination-based defense and our proposed method.
The nucleus in defending LLMs against jailbreak attacks lies in accurately distinguishing harmful (“bad”) input prompts from harmless (“good”) ones, with the precise definition of “good” and “bad” serving as the cornerstone. Existing approaches primarily rely on heuristically defined rules to classify input prompts as harmful or benign (Yi et al., 2024). Prefill-level defenses generally employ pre-defined rules (e.g., keyword lists  (Robey et al., 2023)) or fixed input features (e.g., perplexity thresholds (Alon and Kamfonas, 2023)) to identify malicious prompts. Similarly, response-level defenses leverage post-generation filters, such as harmful content detectors, or pre-established evaluation metrics, including LLM-based evaluations, to flag risky outputs (Wang et al., 2024). However, these static, pre-defined criteria are inherently limited in addressing the nuanced and sophisticated nature of adversarial inputs (Dong et al., 2024). As depicted in Figure 1(a), jailbreak attack prompts manifest in a variety of forms and styles, thereby rendering pre-defined static criteria overly simplistic and rigid. Frameworks that rely on static discrimination between harmful and benign prompts cannot accommodate the inherent complexity and variability of jailbreak attacks, resulting in suboptimal performance in sophisticated real-world scenarios (Huang et al., 2024).
Different from previous approaches relying on static discrimination, we aim to learn dynamic distinction rules that adaptively evaluate the risk of input prompts. Inspired by the concept of Reflective Equilibrium (Rawls, 1971), which achieves consistency through iterative adjustments between judgments and principles, we propose first identifying the input prompts that deviate from the “mirror” and then iteratively adjusting these deviations. The mirror is defined as a dynamically generated prompt that mirrors the syntactic structure of the input while ensuring semantic safety. As illustrated in Figure 1(b), each input prompt is paired with a mirror, and LLMs generate responses for both. If the response to the input significantly deviates from that of its mirror, it is likely to be risky. By incorporating the discrepancy between the responses of the input and its mirror, the discrimination standard becomes dynamically adaptive to each input, as the mirror is generated in a personalized manner. Furthermore, the mirror also serves as a guide for iteratively adjusting the input by minimizing the deviations between the responses of the vanilla input and its mirror.
While mirrors provide a dynamic reference for distinguishing harmful from harmless prompts, their generation and utilization pose significant challenges. First, an effective method is needed to dynamically and accurately generate a suitable mirror that represents the harmless counterpart of the input prompt. Second, a robust metric is essential to quantify the discrepancies between the input and its mirror. Third, an iterative strategy is required to reduce discrepancies between the LLM’s response and the mirror, ultimately ensuring the generation of a safe output.
In this paper, we propose a novel defense method, MirrorGuard, to tackle the aforementioned challenges. Our motivation lies in generating mirrors to serve as dynamic benchmarks for detecting jailbreak attack prompts and guiding the target LLM to produce safe outputs. MirrorGuard consists of three primary modules: the mirror maker, the mirror selector, and the entropy defender. The mirror maker generates desirable mirrors based on the input prompt, implemented as an instruction fine-tuned model. Once the mirrors are generated, the mirror selector identifies the optimal mirror by comparing its syntactic similarity and semantic harmlessness to the input prompt. Finally, the entropy defender iteratively refines the input prompt via mirror comparison, aligning its Relative Input Uncertainty (RIU) value with that of the mirror. Experimental results on popular datasets demonstrate the superiority of our proposal. Our contributions are summarized as follows:
  • To the best of our knowledge, we are the first to investigate the novel problem of dynamic and adaptive defense against jailbreak attacks, based on the innovative concept of “mirror”.
  • We propose MirrorGuard, a novel defense model that dynamically generates mirrors to adaptively detect attacks. An iterative alignment strategy is proposed to minimize deviations between inputs and their mirrors.
  • We conduct extensive experiments on popular benchmarks and various jailbreak attack methods, demonstrating that our approach outperforms baselines in reducing attack success rates while maintaining overall performance.
Figure 2: The overview of the proposed MirrorGuard model, including the mirror maker, the mirror selector, and the entropy defender via mirror comparison.
2 Preliminary
Entropy.
Entropy is a foundational concept in information theory (Pimentel et al., 2021). As a measure of uncertainty in a LLM’s output, entropy reflects the model’s confidence of its information content (Araabi et al., 2024). Given a random variable $X$ with the set $[x_{0},\dots,x_{n}]$ of possible outcomes, the Shannon entropy (Shannon, 1948) is defined as follows:
$H(X)=-\sum_{i}{P(x_{i})\log{P(x_{i})}},$ (1)
where $P(x_{i})\geq 0$, and $\sum_{i=0}^{n}P(x_{i})=1$.
Attention Entropy.
In transformer-based LLMs, encoder applies scaled-dot product self-attention over the input tokens to compute $N$ independent attention heads. Let $E=[\boldsymbol{e_{0}},\dots,\boldsymbol{e_{d_{s}}}]$ be the sequence of input embeddings, with $\boldsymbol{e_{i}}\in\mathbb{R}^{d_{m}}$. For the $h$-th attention head and $i$-th token position, each embedding $\boldsymbol{e_{i}}$ is projected into a query $\boldsymbol{q}_{h,i}$, a key $\boldsymbol{k}_{h,i}$, and a value $\boldsymbol{v}_{h,i}$. The attention weight for the $i$-th token can be calculated as follows:
$\alpha_{h,i}=\operatorname{softmax}\left(\frac{\boldsymbol{q_{h,i}}^{T}K_{h}}{\sqrt{d_{k}}}\right),$ (2)
where $K_{h}$ is the matrix of keys for the $h$-th head, and $d_{k}$ is the dimensionality of the key vectors.
Let $\alpha_{h,i,j}$ denote the $j$-th entry of $\alpha_{h,i}$, which quantifies the importance of the $j$-th token in determining the representation of the $i$-th token for the $h$-th attention head. These weights satisfy $\alpha_{h,i,j}\geq 0$ and $\sum_{j=0}^{d_{s}}\alpha_{h,i,j}=1$. Following previous work (Ghader and Monz, 2017), we compute the entropy in the attention heads by interpreting each token’s attention distribution as a probability mass function of a discrete random variable. In this context, the input tokens (or their embeddings) are treated as the possible outcomes, and the attention weights serve as their probabilities. To analyze the collective contribution of all attention heads, the attention weights are first averaged across heads to compute the mean attention as follows:
$\bar{\alpha}_{i,j}=\frac{1}{N}\sum_{h=1}^{N}\alpha_{h,i,j},$ (3)
where $\bar{\alpha}_{i,j}$ represents the mean probability that the $j$-th token contributes to the representation of the $i$-th token across all heads. The attention entropy for the $i$-th token can be calculated as follows:
$H_{i}=-\sum_{j=0}^{d_{s}}\bar{\alpha}_{i,j}\log\bar{\alpha}_{i,j}.$ (4)
Attention entropy measures the degree of contextualization in constructing model’s higher-level embeddings, where higher value of entropy indicates greater uncertainty (Attanasio et al., 2022).
3 Methodology
Figure 2 (a) illustrates the joint framework of the MirrorGuard paradigm. Given a input jailbreak attack prompt $p$, MirrorGuard firstly generates a set of candidate mirrors {$l_{1}$, $l_{2}$,…, $l_{n}$} via the Mirror Maker. Then, Mirror Selector selects the most satisfied mirror $l_{i}^{\prime}$ which has the same linguistic characters with the input $p$. $l_{i}^{\prime}$ is expected to compare the behavioral inconsistencies in the sentence understanding of the LLM. Finally, Entropy Defender via Mirror Comparison calculates RIU between current input and mirrors to identify the jailbreak attack prompt, and guides the LLM toward generating more confident and reliable responses through the multiple-query guidance process.
3.1 Mirror Maker
The objective of mirror maker is to produce mirrors that share the similar linguistic features with the input jailbreak attack prompt. An effective mirror is expected to satisfy three constraints. (1) Length constraint: the mirror is expected to have a token count similar to that of the input prompt, allowing for a fair comparison between prompt pairs (Steindl et al., 2024). (2) Syntax Constraint: the mirror needs to follow syntactic rules consistent with the input prompt to control variables in the LLM’s comprehension of the input (Zhu et al., 2024). (3) Sentiment constraint: the mirror is also intended to have a non-negative sensitive polarity, serving as a safe and standard reference (Huang et al., 2024).
Constrained Instruction Tuning. In light of the stochastic nature of LLM’s outputs and their reliance on general-purpose training, it is intractable for existing models to directly generate mirrors due to the lack of explicit control mechanisms. Drawing inspiration from prior work (Zhou et al., 2023), we employ fine-tuning guided by natural language instructions and demonstrations to construct an instruction-tuned model.
The fine-tuning process begins by constructing a training dataset consisting of constraint–text pairs, where constraints are verbalized into natural language and paired with sampled texts. As displayed in Figure 2 (b), three types of constraints are considered: length constraint, syntax constraint, and sentiment constraint. For the length constraint, we define an interval based on a configurable parameter $\lambda$, requiring the output length to fall between $\lambda n$ and $\lambda(n+1)$ words, where $n$ is an integer. For the syntax constraint, syntactic structures are enforced through linearized syntactic parse trees, such as (S (NP (PRP *)) (VP (VBD *) (NP (DT *) (NN *)))), where the asterisk (*) acts as a wildcard, allowing flexible word choice while preserving the specified syntactic structure. As for the sentiment constraint, GPT-4o is employed to assign sentiment labels (e.g., positive, neutral, negative) to randomly sampled sentences, ensuring the dataset contains diverse sentiment-conditioned examples.
After the process of data synthesis, the model is fine-tuned by using a combination of instruction tuning (Wei et al., 2022) and meta-in-context learning (Min et al., 2022). For each training example, the natural language instructions for the constraints are prepended with 5 demonstrations, which consist of constraint–output pairs. These demonstrations are either of the same constraint type or a composition of multiple constraints. The model is then fine-tuned by using maximum likelihood estimation and teacher forcing technology to ensure training stability and constraint compliance (Zhou et al., 2023).
Mirror Generation. The mirror generation focuses on producing mirror $l$ on basis of the input prompt $p$ and the instruction-tuned model $\mathcal{G}$, formalizing as $l=\mathcal{G}(p)$. This process begins by concatenating the instruction prompt with the input prompt, which is then fed into the instruction-tuned model to generate candidate mirrors. To ensure the instruction-tuned model can simultaneously handle multiple constraints, each constraint is individually described and subsequently combined using the conjunctive “and”. For instance, given the sentence “He makes a cake” with constraints on both length and syntactic structure, the final instruction would be formulated as: “Write something that has 1 to 5 words and follows the part-of-speech tag sequence PRP VERB DET NOUN”. Once the input prompt and constraint are sent to the instruction-tuned model, $k$ responses will be generated as a set of candidate mirrors.
3.2 Mirror Selector
Owing to the inherent stochasticity in the generation process of LLMs, not all candidate mirrors are capable of adhering to these constraints. To address this challenge, a mirror selector is proposed to identify the most suitable mirrors based on three predefined criteria: length consistency, syntax consistency, and sentiment consistency. Specifically, length consistency ensures that the selected mirror closely matches the original input in terms of token count, while syntax consistency requires the mirror to preserve a similar grammatical structure. Sentiment consistency evaluates whether the polarity of the mirror remains non-negative.
To evaluate and filter the candidate mirrors, an LLM-based classifier (e.g., GPT-4o) is employed. This classifier assesses each candidate mirror against the three constraints and assigns judgment tags, such as “true” or “false” to indicate whether the criteria are satisfied. Only the mirrors that simultaneously satisfy all three criteria are considered suitable. Among these suitable mirrors, the first five are selected as the final choices to ensure robustness and effectiveness. The details of mirror selector process are provided in Appendix B.
3.3 Entropy Defender via Mirror Comparison
(a) $l_{jail}$ vs $l_{mirror_{1}}$
(b) $l_{mirror_{1}}$ vs $l_{mirror_{2}}$
Figure 3: Comparison of attention entropy with jailbreak attack prompts and harmless prompts.
Following the mirror maker and mirror selector, a set of desirable mirrors is generated. However, the integration of these mirrors for effective defense remains obscure. We further propose the entropy defender to leverage the discrepancies between input prompts and their mirrors to dynamically assess and mitigate the risks.
3.3.1 Discrepancy Quantification via RIU
A straightforward approach to analyze the divergence between an input prompt and its mirror might involve directly comparing their outputs. However, this comparison is problematic, as substantial semantic differences between the input and mirror can render the outputs incomparable. Prior research shows that jailbreak attacks induce inconsistent model responses, indicating that uncertainty can effectively quantify such discrepancies (Steindl et al., 2024). Building on this, we propose using uncertainty to quantify the differences in the model’s inference of the input and its mirror.
Relative Input Uncertainty (RIU). Uncertainty in LLMs can be effectively quantified using entropy, a measure that captures the unpredictability of the internal states (Araabi et al., 2024). Specifically, we employ attention entropy in Equation (4) to provide a robust assessment of the model’s internal uncertainty regarding its processing of different input components. To measure the divergence between an input prompt and its mirrored counterpart, we further adopt Information Gain (IG)  (Hu et al., 2024). This metric offers a dynamic and precise means of quantifying uncertainty, capturing the variations in entropy between two states. In our context, the entropy of the mirror scenario represents uncertainty under idealized conditions, while the entropy of the current input reflects the actual uncertainty encountered by the model. Formally, for the $i$-th token, IG is defined as follows:
$|IG_{current}|=\frac{1}{d_{s}}\sum_{i=1}^{d_{s}}\big{|}H_{l_{input}}^{i}-H_{l_{1}}^{i}\big{|},$ (5)
where $d_{s}$ denotes the total number of tokens in the input sequence, $H_{input}^{i}$ and $H_{l_{1}}^{i}$ represent the token-level attention entropy for the current input and the mirror, $mirror_{1}$. While IG quantifies the absolute divergence between an input prompt and its mirror, it lacks a reference standard to contextualize the magnitude of this difference. To address this, we propose Relative Input Uncertainty (RIU) to measure the relative divergence in attention entropy between the input prompt and its mirror. RIU compares two scenarios: <$mirror_{1}$ ($l_{1}$), $mirror_{2}$ ($l_{2}$)> and <$mirror_{1}$ ($l_{1}$), input prompt ($l_{input}$)>. Let $|IG_{reference}|$ serve as the expected reference standard to assess the divergence in the first scenario, the RIU can be formulated as the follows:
$RIU=\frac{|IG_{reference}|}{|IG_{current}|}=\frac{\big{|}H_{l_{1}}^{i}-H_{l_{2}}^{i}\big{|}}{\big{|}H_{l_{input}}^{i}-H_{l_{1}}^{i}\big{|}},$ (6)
where $H_{l_{2}}^{i}$ represents the token-level attention entropy for $mirror_{2}$ at the $i$-th token.
Preliminary Analysis of RIU. To validate the effectiveness of RIU in quantifying discrepancies, we conduct experiments on the AdvBench dataset (Zou et al., 2023) to analyze the RIU values for different types of inputs. Several popular jailbreak attack methods are selected to compare their impact, including PAIR (Chao et al., 2023), TAP (Mehrotra et al., 2023), DrAttack (Li et al., 2024), and BaitAttack (Pu et al., 2024). Four open-source LLMs are used to explore the effectiveness of RIU, including Llama2-7b-chat (Touvron et al., 2023), Llama3-8b, Vicuna-7b (Chiang et al., 2023), and Mistral-7b (Jiang et al., 2023).
(a) Llama2-7b
(b) Llama3-8b
(c) Vicuna-7b
(d) Mistral-7b
Figure 4: The comparison of RIU under different attack methods across four LLMs.
Figure 3 compares the differences in $IG$ between $\langle l_{jail},l_{mirror_{1}}\rangle$ and $\langle l_{mirror_{1}},l_{mirror_{2}}\rangle$ under the Llama2-7b-chat model, where $l_{jail}$ refers to a jailbreak attack prompt from PAIR and $l_{mirror_{1}}$ and $l_{mirror_{2}}$ denote its mirrors. In Figure 3 (a), the differences of attention entropy between $l_{jail}$ and $l_{mirror_{1}}$ exhibit sharp peaks, reflecting a volatile divergence in attention distribution. In contrast, Figure 3(b) shows subdued low peaks for $l_{mirror_{1}}$ and $l_{mirror_{2}}$, indicating smoother, more stable variations. These observations reveal that the model’s attention distribution is highly uneven and inconsistent with jailbreak attack prompts. While when handling harmless prompts that share similar syntax but differ in semantics, the attention distribution exhibits greater stability and consistency.
To further analyze RIU, we visualize the RIU differences across multiple models in Figure 4. The sub-figures show that for mirror pairs, the values of RIU are consistently close to 1.0. For the comparisons between jailbreak prompts and mirrors, the value of RIU differ significantly and are much lower than 1.0. These findings show that RIU can effectively quantify the discrepancies between input and its mirror, which provides a robust foundation for detecting and defending jailbreak attacks.
3.3.2 RIU-based Defense
The RIU-based defense aims to dynamically measure the risk of an input prompt by refining the input prompt based on the RIU scores. As illustrated in Figure 2 (d), let $p$ denote the input prompt, $l_{1}$ and $l_{2}$ represent two semantically distinct mirrors corresponding to $p$, the current $IG$ between $P$ and $l_{1}$ can be computed as follows:
$|IG_{current}|=|H(p)-H(l_{1})|,$ (7)
where $H(p)$ and $H(l_{1})$ are the attention entropy of the input prompt and its mirror, respectively. Similarly, the target $IG$ can be calculated as follows:
$|IG_{target}|=|H(l_{1})-H(l_{2})|.$ (8)
Then, the $RIU_{attention}$ is formulated as:
$|RIU_{attention}|=\frac{|IG_{target}|}{|IG_{current}|}.$ (9)
From Equation (6), when the input prompt is benign, the numerator and denominator tend to be similar (i.e., $RIU_{attention}$ close to 1.0), as both quantify the deviations between two safe yet syntactically similar sentences. If $RIU_{attention}$ exceeds a predefined threshold $\sigma$ (close to 1.0), then the LLM is classified as harmless, and the model proceeds with response generation. However, if $RIU_{attention}$ falls below the threshold, the input is flagged as potentially harmful, triggering a multiple-query guidance process. This process iteratively refines the input prompt through strategies such as simplification (Steindl et al., 2024)(e.g., “Please simplify the current sentence”) and multiple queries (Hu et al., 2024) until $RIU_{attention}$ exceeds the threshold. The algorithm for MirrorGuard is detailed in Appendix C.
4 Experiment
4.1 Experimental Settings
Datasets. Following previous work Xu et al. (2024), two harmful query datasets, Advbench and HEx-PHI, are adopted to evaluate the effectiveness of various defense methods (Qi et al., 2023). Advbench contains 520 malicious queries (Chao et al., 2023), while HEx-PHI consists of 330 offensive questions. To assess the performance of LLMs on harmless prompts, several datasets, such as AlpacaEval (Dubois et al., 2023) and VicunaEval (Chiang et al., 2023) are adopted.
Target LLMs. The defense method is evaluated on two widely used open-source LLMs: Llama2-7b-chat (Touvron et al., 2023) and Vicuna-7b (Chiang et al., 2023). Following previous work (Zhou et al., 2023), T0-11b is utilized as the base model of instruction-tuned model because it has previously been instruction-tuned on many NLP tasks.
Attacks. To address the effectiveness of MirrorGuard, six representative attacks are selected to be compared: GCG (Zou et al., 2023), AutoDAN (Liu et al., 2024), PAIR (Chao et al., 2023), TAP (Mehrotra et al., 2023), DrAttack (Li et al., 2024) and BaitAttack (Pu et al., 2024).
Baselines. Following the previous work (Xu et al., 2024), five SOTA defense mechanisms are considered as baselines, including detection-based (Perplexity Filter (Alon and Kamfonas, 2023)) and mitigation-based methods (ICD (Wei et al., 2023), Paraphrase (Jain et al., 2023), Self-Reminder (Xie et al., 2023) and SafeDecoding (Xu et al., 2024)).
Model Defense Harmful Benchmark($\downarrow$) Jailbreak Attacks($\downarrow$) Average
Optimization Generation Indirect
AdvBench HEx-PHI GCG AutoDAN PAIR TAP DrAttack BaitAttack
Llama2 No Defense 0.000 0.002 0.382 0.575 0.174 0.187 0.511 0.654 0.414
Perplexity Filter 0.000 0.002 0.000 0.575 0.174 0.187 0.511 0.654 0.350
ICD 0.000 0.000 0.000 0.000 0.147 0.156 0.231 0.255 0.132
Paraphrase 0.002 0.003 0.004 0.000 0.124 0.151 0.275 0.298 0.142
Self-Reminder 0.000 0.000 0.000 0.000 0.139 0.147 0.301 0.342 0.155
SafeDecoding 0.000 0.001 0.004 0.000 0.004 0.005 0.009 0.011 0.006
MirrorGuard(Ours) 0.000 0.001 0.000 0.000 0.002 0.002 0.006 0.009 0.003
Vicuna No defense 0.008 0.172 0.910 0.847 0.815 0.823 0.854 0.979 0.871
Perplexity Filter 0.008 0.154 0.000 0.847 0.815 0.823 0.854 0.979 0.720
ICD 0.000 0.006 0.711 0.805 0.537 0.544 0.832 0.891 0.720
Paraphrase 0.143 0.237 0.225 0.714 0.679 0.680 0.811 0.887 0.666
Self-Reminder 0.000 0.008 0.428 0.702 0.485 0.486 0.825 0.854 0.630
SafeDecoding 0.000 0.001 0.000 0.000 0.008 0.009 0.013 0.025 0.009
MirrorGuard (Ours) 0.000 0.001 0.000 0.000 0.002 0.007 0.010 0.012 0.005
Table 1: The ASR results of different LLMs under various defense methods. The best results are highlighted in bold.
Defense Llama2 Vicuna
Perplexity Filter 0.982 $\times$ 0.984 $\times$
ICD 1.013 $\times$ 1.014 $\times$
Paraphrase 2.148 $\times$ 1.796 $\times$
Self-Reminder 1.011 $\times$ 1.012 $\times$
SafeDecoding 1.033 $\times$ 1.072 $\times$
MirrorGuard(Ours) 1.043 $\times$ 1.075 $\times$
Table 2: This table summarizes ATGR of MirrorGuard and the baseline defense approaches.
Defense AlpacaEval($\uparrow$) VicunaEval($\uparrow$)
WinRate Rouge-L WinRate Rouge-L
No Defense 69.6% 0.453 92.5% 0.541
Perplexity Filter 61.7% 0.306 70.3% 0.392
ICD 62.3% 0.381 75.0% 0.456
Paraphrase 34.1% 0.257 43.6% 0.227
Self-Reminder 58.3% 0.289 65.0% 0.316
Self-Decoding 66.3% 0.439 90.4% 0.525
MirrorGuard(Ours) 68.6% 0.443 91.2% 0.535
Table 3: Impact of defenses on LLMs’ general performance on benign datasets.
Metrics. Following previous work Xu et al. (2024), the attack success rate (ASR) is used to measure the effectiveness of the defense methods, which is defined as: $ASR=\frac{\#\text{successful prompts}}{\#\text{total prompts}}$. The success of a jailbreak attack is evaluated by GPT-4o (Qi et al., 2023). The Average Token Generation Time Ratio (ATGR) is used to assess the time cost of all defense methods (Xu et al., 2024). It is defined as $ATGR=\frac{T_{\text{defense}}}{T_{\text{non-defense}}}$, where $T_{\text{defense}}$ indicates the average token generation time when defenses are implemented, and $T_{\text{non-defense}}$ indicates the average token generation time when defenses are not implemented. Moreover, the WinRate and Rouge-L scores are used to evaluate the general performance of LLMs in dealing with harmless tasks (Jiang et al., 2024).
Implementation Details. The details of implementation settings are given in Appendix B.
4.2 Evaluation of Defense Effectiveness
Table 1 compares the results of ASR for evaluating the effectiveness of RIU and baselines against six jailbreak attacks. One can see that for models with weak safety alignment, such as Vicuna, MirrorGuard significantly reduces ASR, outperforming almost all baseline defenses. For instance, while most other defenses fail to mitigate indirect jailbreaks, MirrorGuard succeed to achieve the ASR close to 0. While for models that are well aligned, such as Llama2, MirrorGuard reduces the ASR of all attacks to nearly 0. This can contribute to the dynamic generated mirrors in MirrorGuard, which is adaptively tailored to construct syntactic structures that align closely with the input.
4.3 Evaluation of Defense Efficiency
In Table 2, we present a comparison of the ATGR with and without the implementation of defense mechanisms. The value of ATGR under MirrorGuard is 1.043$\times$ for Llama2 and 1.075$\times$ for Vicuna, demonstrating a small computational overhead and maintaining efficiency comparable to the baseline methods. When compared to approaches like Perplexity Filter (0.982$\times$ for Llama2 and 0.984$\times$ for Vicuna) and ICD (1.013$\times$ and 1.014$\times$, respectively), MirrorGuard introduces a slightly higher overhead. This is due to its more sophisticated mechanisms for enhancing the robustness of LLMs against jailbreak attack prompts. However, this increase in computational cost is negligible since wll the value of ATGR under MirrorGuard is close to 1.000. Overall, these results affirm that the slight computational trade-offs associated with MirrorGuard are well-justified.
Target LLMs Llama2 Vicuna
Attack Methods GCG AutoDAN PAIR TAP DrAttack BaitAttack GCG AutoDAN PAIR TAP DrAttack BaitAttack
No defense 0.382 0.575 0.174 0.187 0.511 0.654 0.910 0.847 0.815 0.823 0.854 0.979
w/o length 0.382 0.575 0.174 0.187 0.511 0.654 0.910 0.847 0.815 0.823 0.854 0.979
w/o syntax 0.185 0.272 0.081 0.098 0.186 0.243 0.562 0.447 0.539 0.516 0.416 0.478
w/o sentiment 0.033 0.021 0.035 0.040 0.069 0.071 0.048 0.037 0.054 0.055 0.103 0.119
w/o multiple-query 0.000 0.000 0.012 0.016 0.018 0.020 0.000 0.000 0.018 0.020 0.074 0.081
MirrorGuard 0.000 0.000 0.002 0.002 0.006 0.009 0.000 0.000 0.002 0.007 0.010 0.012
Table 4: Ablation study on the constraints of mirror maker and the multiple-query guidance in entropy defender.
4.4 Evaluation of General Performance
Despite enhancing the safety of LLMs, ensuring the helpfulness of LLMs is also important. Table 3 summarizes the general performance results of Llama2 in dealing with benign tasks under various defense methods. MirrorGuard has the least impact on the LLM’s general performance. Specifically, it achieves a WinRate of 68.6% on AlpacaEval and 91.2% on VicunaEval. And the Rouge-L scores closely match those of undefended Llama2. This is due to that the introduction of mirrors can effectively distinguish between benign prompts and jailbreak attack prompts. Moreover, the multiple queries of MirrorGuard can help the model understand the input prompt.
4.5 Ablation Study
Table 4 presents the ASR results of Llama2 and Vicuna without the constraints of mirror maker and the multiple-query guidance in entropy defender.
Template Constraint. From Table 4, it can be observed that the length constraint is the most crucial. Removing it causes performance to drop nearly to zero, as the RIU computation depends on sentences having the same number of tokens for accurate attention. The syntax constraint is also important, with its removal resulting in a performance decline of over 10%, as differing syntactic structures introduce variability that hinders the RIU computation. In contrast, the sentiment constraint has the least impact, as mirror sentences typically do not generate harmful content without explicit guidance. Therefore, removing this constraint introduces less variability than the other two constraints.
Multiple-Query Guidance. Table 4 indicates that removing the multiple-query guidance can lead to only a slight degradation of performance, which suggests that a single round of querying can also produce effective results. This is due to the targeted and adaptive nature of mirror, which helps the model effectively distinguish jailbreak attack prompts and harmless input prompts. And the multiple-query guidance serves as a complementary component, it enhances the robustness of LLMs against more complex jailbreak attacks.
4.6 Hyper-parameter Sensitivity Analysis
0011112222333344445555666677770.000.000.000.000.200.200.200.200.400.400.400.400.600.600.600.60number of candidate mirrorsASR
(a) The trend of ASR with the increasing number of candidate mirrors.
0.600.600.600.600.650.650.650.650.700.700.700.700.750.750.750.750.800.800.800.800.850.850.850.850.900.900.900.900.500.500.500.500.600.600.600.600.700.700.700.700.800.800.800.800.900.900.900.901.001.001.001.00the value of $\sigma$Accuracy
(b) The trend of accuracy with the increasing value of threshold $\sigma$.
Figure 5: Hyperparameter sensitivity analysis.
Here we conduct hyper-parameter sensitivity analysis on the two parameters: the number of candidate mirrors, and the threshold of RIU.
The number of candidate mirrors. Based on experiments using the BaitAttack method under the Llama2 model, Figure 5(a) shows the impact of varying the number of candidate mirrors on ASR. As the number of candidates increases, the ASR steadily drops. When the number reaches 4, ASR falls below 0.3, marking a substantial improvement. When 5 or more candidates are used, ASR approaches zero and remains low. These findings suggest that 5 candidate mirrors offer an optimal trade-off between defense effect and efficiency.
The threshold of RIU. Figure 5(b) shows the trend of accuracy in distinguishing between harmful and harmless prompts as the value of RIU’s threshold $\sigma$ increases. The dataset consists of 100 benign and 100 jailbreak attack prompts, and the adopted model is Llama2. The accuracy shows a clear increase as the $\sigma$ rises from 0.60 to 0.80, reaching a peak of 0.97 when $\sigma$ is 0.80. However, beyond this point, accuracy starts to decline, with a noticeable drop to 0.75 when $\sigma$ is 0.90. Thus, the optimal threshold $\sigma$ is around 0.80.
5 Conclusion
To overcome the limitations of static methods in detecting harmful inputs, we propose MirrorGuard, a dynamic defense strategy using “mirrors” to differentiate safe from malicious prompts. We introduce RIU, a novel metric to quantify discrepancies between an input and its mirror. Guided by RIU, MirrorGuard ensures reliable responses while preserving LLMs’ general reasoning abilities. Evaluations demonstrate its superior performance in detecting and mitigating jailbreak attacks.
6 Limitation
Based on the relation between inputs and their entropy difference, mirror is proposed to reflect the distinction of the inputs with jailbreak or harmless elements. In addition, a novel defense paradigm which is called as MirrorGuard is put forward to analyze the difference between current prompt and mirror with accurately ability in detecting jailbreak prompts. Some experiments are given to prove the effectiveness of mirror and MirrorGuard. However, there are still some limitations for mirror and MirrorGuard as follows: (1) Mirror is only used as a baseline for comparing attention entropy difference. While effective for detecting the anomalies of input, it may overlook subtle adversarial manipulations beyond attention patterns. Future work can explore more comprehensive metrics to better understand and address such complexities. (2) The validity of MirrorGuard is only demonstrated across four open-source models under six common types of jailbreak attacks. The generality of this method is still to be further proved.
Appendix A Related Work
Existing defensive strategy can be mainly divided into two types: input-based and output-based defending strategies (Zeng et al., 2024).
A.1 Prefill-level defense
In prefill-level defense, the inputs will be optimized to induce LLMs to reject harmful questions. The optimizing techniques include adding prompts with safety system or filtering the input. Adding prompt refers to delving into the intent of input before decoding (Zhang et al., 2024). Perplexity filtering is often used to detect the adversarial suffixes as the harmful input signal before generating a response (Alon and Kamfonas, 2023). The input of the encoder can be modified to defend the jailbreak attack (Zhao et al., 2024). However, modifying input for jailbreak attacks is a deficient consumption and efficient behavior. Hence, many improved methods have been successfully proposed to defend jailbreak attacks from user input, such as GCG (Zou et al., 2023), AutoDAN (Zhu et al., 2023), Evil Geniuses (Tian et al., 2023). Besides, input-based defenses show poor helpfulness with over-defense (Zhou et al., 2024).
A.2 Response-level defense
Response-level defenses can be used to enhance the security of LLMs by judging the generated response (Zeng et al., 2024). For instance, Self-Examination can realize checking the response itself by using a pre-defined prompt (Phute et al., 2023). Self-Decoding can capture the safety disclaimers from top tokens and amplify their token probabilities (Xu et al., 2024). Whereas, response level defenses must generate all the outputs before judging, which affects the efficiency of LLMs. While the attention mechanism is introduced in MirrorGuard monitors step-by-step, which can effectively reduce the time in realizing safety.
Appendix B Implementation Details
B.1 Training Details
Building on prior work by INSTRUCTCTG (Zhou et al., 2023), we generate 10 thousand constraint-text training pairs for each constraint category (length, syntax, and sentiment), along with 2000 pairs for both the development and test sets. The data is all randomly sampled from the C4 dataset (Raffel et al., 2019). And the model is fine-tuned with a learning rate of 5e-5.
B.2 Constraint Template Samples
As shown in Table 5, for each constraint type, natural language templates are designed to verbalize the constraints into natural language prompts. The template is the same as INSTRUCTCTG (Zhou et al., 2023).
LENGTH TARGET: I am going to finish the 3rd season of My Brilliant Friend tonight.
CONSTRAINTS: length of 10-15 words
INSTRUCTION-1: Write something that has 10 to 15 words.
INSTRUCTION-2: Write something with 10 to 15 words.
SYNTAX TARGET: Avatar is a science fiction film directed by James Cameron.
CONSTRAINTS: NOUN VERB DET NOUN NOUN NOUN VERB PREP NOUN NOUN PUNCT
INSTRUCTION-1: Write something with a part-of-speech sequence NOUN VERB DET NOUN NOUN NOUN VERB PREP NOUN NOUN PUNCT.
INSTRUCTION-2: Write something that results in the part-of-speech sequence NOUN VERB DET NOUN NOUN NOUN VERB PREP NOUN NOUN PUNCT after part-of-speech tagging with Spacy.
SENTIMENT TARGET: Love this place. Never had a bad meal. Good portions and great people.
CONSTRAINTS: the sentiment is positive
INSTRUCTION-1: Write something with positive sentiment.
INSTRUCTION-2: Write something positive.
Table 5: Example texts, constraints, and the corresponding verbalized instructions. The instructions are highlighted in red. Note that there are multiple possible ways to verbalize instructions for each task.
B.3 Selection Template
You will act as a classifier to evaluate a given candidate sentence against three specific criteria: Length Consistency, Syntax Consistency, and Sentiment Consistency. For each criterion, you must determine whether the sentence satisfies the requirement and respond with either “True” or “False”. The evaluation must be based on the following criteria:
  • Length Consistency: The number of tokens in the candidate sentence must match the token count of the original input prompt.
  • Syntax Consistency:The candidate sentence must adhere to the grammatical structure specified by the provided syntax constraint. This could include maintaining a prescribed parse tree structure or following a specific part-of-speech tagging sequence. For example, if the given syntax constraint specifies a structure such as (S (NP (DT *) (NN *)) (VP (VBD *) (NP (DT *) (JJ *) (NN *))), the candidate must conform to this representation.
  • Sentiment Consistency: The candidate sentence must exhibit positive or neutral sentiment without any harmful, offensive, or ambiguous language. Sentences with clearly negative sentiment or sentences that fail to maintain a safe and neutral tone are considered invalid.
You will receive the following input information:
  • Input Prompt: “<original input prompt>”
  • Candidate Sentence: “<candidate mirror>”
  • Length Constraint: “<expected token count>”
  • Syntax Constraint: “<required syntactic structure or parse tree>”
The Output Format:
  • Length Consistency: True/False
  • Syntax Consistency: True/False
  • Sentiment Consistency: True/False
B.4 More Details fo Metrics
The WinRate and Rouge-L scores are used to evaluate the general performance of LLMs in dealing with harmless tasks (Jiang et al., 2024). This evaluation of WinRate uses 100 queries from AlpacaEval and 80 queries from VicunaEval. And GPT-4o is used as the evaluator. Additionally, the two metrics are both measured by using the generated responses of text-davinci-003 as reference points.
Appendix C Algorithm
Algorithm 1 MirrorGuard Algorithm
1:Prompt $p$, Mirror Maker $\mathcal{G}$, Threshold $\epsilon$
2:Safe and reliable output
3:Step 1: Mirror Maker
4:Generate mirrors: $\{l_{1},l_{2},\dots,l_{n}\}\leftarrow\mathcal{G}(p)$
5:Step 2: Mirror Selector
6:Select mirrors $\{l^{\prime}_{1},l^{\prime}_{2}\}$ that satisfy:
7:    Length, syntax, and sentiment constraints
8:Step 3: Entropy Defender via Mirror Comparison
9:Compute |$IG_{\text{current}}|\leftarrow\text{IG}(p,l^{\prime}_{1})$
10:Compute |$IG_{\text{target}}|\leftarrow\text{IG}(l^{\prime}_{1},l^{\prime}_{2})$
11:while $|RIU_{\text{attention}}|>\epsilon$ do
12:     Update prompt: $p^{\prime}\leftarrow\text{Simplify/Modify}(p)$
13:     Recompute |$IG_{\text{current}}|\leftarrow\text{IG}(p^{\prime},l^{\prime}_{1})$
14:end while
15:Output: final response